OpenStudy (anonymous):

A car travels at 15 m/s for 10 s. It then speeds up with a constant acceleration of 2.0 m/s2 for 15 s. At the end of this time, what is the magnitude of its velocity? HELP

OpenStudy (anonymous):

answers are : a. 45 m/s b. 15 m/s c. 30 m/s d. 375 m/s

OpenStudy (anonymous):

V=V(0)+at so we get V=15m/s+2/m^2 15s=45m/s so a is the right answer

OpenStudy (anonymous):

Thank you